OpenStudy (cwrw238):

Please try this question from a math test. I am baffled by it. 4fccab66e4b0c6963ad793df-smishra-1338813325025-copyofnewmicrosoftofficeworddocument.docx

OpenStudy (anonymous):

i dont think it uploaded properly try again and ill try and help

OpenStudy (cwrw238):

i'll draw it

OpenStudy (anonymous):

ok :)

OpenStudy (cwrw238):

If a, b, c are positive and are pth, qth and rth terms of a GP; then show without expanding that | loga p 1 | | logb q 1 | = 0 | logc r 1 |

OpenStudy (anonymous):

sorry i didnt read

OpenStudy (cwrw238):

i think its a determinant

OpenStudy (anonymous):

what does GP stand for?

OpenStudy (cwrw238):

geometric series

OpenStudy (anonymous):

ah nice. ok give me a minute

OpenStudy (cwrw238):

i assume that they are consecutive terms - can we do that?

OpenStudy (anonymous):

unfortunately not, but we do know their positions, p,q,r

OpenStudy (anonymous):

the nth term of a GP is A(r^(n-1)) where A is the first term and r is the common ratio

OpenStudy (anonymous):

to avoid confusion lets use d instead of r

OpenStudy (anonymous):

so lets call the first term of the progression "f" and the common ratio "d"

OpenStudy (anonymous):

can you see this: \[a= fd^{p-1}\]

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

cool have a go playing with that, and see how far you can get it

OpenStudy (cwrw238):

b = fd^(q-1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that was a lovely problem :) made me very happy. i haven't really come across logarithms intertwined with matrices before. thanks for showing me

OpenStudy (anonymous):

hows it going?

OpenStudy (cwrw238):

i'm struggling - whats confusing me is that it says do not expand

OpenStudy (anonymous):

ah ok well ill put up my solution and you can read through as far as you like

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

\[\left|\begin{matrix}log(a) & p & 1\\ log(b) & q & 1\\ log(c) & r & 1\end{matrix}\right| \]

OpenStudy (anonymous):

\[\left|\begin{matrix}\log(fd^{p-1}) & p & 1\\ \log(fd^{q-1}) & q & 1\\ \log(fd^{r-1}) & r & 1\end{matrix}\right|\]

OpenStudy (anonymous):

now just to check, are you familiar with how you can manipulate determinants, eg row operations?

OpenStudy (cwrw238):

yes - i got that far! - no more!

OpenStudy (cwrw238):

only vaguely

OpenStudy (cwrw238):

thats why I'm stuck

OpenStudy (anonymous):

lets start by applying some log laws first we break apart the logarithms like this: \[\left|\begin{matrix}\log(d^{p-1}) + \log(f) & p & 1\\ \log(d^{q-1}) + \log(f) & q & 1\\ \log(d^{r-1}) + \log(f) & r & 1\end{matrix}\right|\]

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

then: \[\left|\begin{matrix}\ (p-1)\log(d) + \log(f) & p & 1\\ \ (q-1)\log(d) + \log(f) & q & 1\\ \ (r-1)\log(d) + \log(f) & r & 1\end{matrix}\right|\]

OpenStudy (cwrw238):

yea i follow that

OpenStudy (anonymous):

ok, now we do some row operations. remember row operations do not change the value of the determinant so we can use them here. first we do R1 - R2 ( that's row 1 - row 2) you should get this: \[\left|\begin{matrix}\ (p-q)\log(d) & p-q & 0\\ \ (q-1)\log(d) + \log(f) & q & 1\\ \ (r-1)\log(d) + \log(f) & r & 1\end{matrix}\right|\]

OpenStudy (anonymous):

now we do R2- R3 : \[\left|\begin{matrix}\ (p-q)\log(d) & p-q & 0\\ \ (q-r)\log(d) & q-r & 0\\ \ (r-1)\log(d) + \log(f) & r & 1\end{matrix}\right|\]

OpenStudy (cwrw238):

oh yes i remember doing something like this - to find inverse of matrix i think

OpenStudy (cwrw238):

i think i know whats coming - when you calculate the determinant the zeros will be significant

OpenStudy (anonymous):

probably because you need determinants to find inverses. anyway back to the determinant :D now can you see IF we were to expand the determinant using column 3 to expand, the two left elements of R3 would not contribute to the determinant? yeah you guessed it, but we are not allowed to expand, we'll have to keep manipulating.

OpenStudy (anonymous):

because of the two existing zeros we can do this: \[\left|\begin{matrix}\ (p-q)\log(d) & p-q & 0\\ \ (q-r)\log(d) & q-r & 0\\ \ 0 & 0 & 1\end{matrix}\right|\] do you understand or should i explain this step more?

OpenStudy (cwrw238):

to be honest no

OpenStudy (anonymous):

ok no worries, ill explain :)

OpenStudy (cwrw238):

yes i think i see now - those elements in row 3 are multiplied by 0

OpenStudy (anonymous):

how about trying to do the operation (column 2) - r(column 3) here: \[\left|\begin{matrix}\ (p-q)\log(d) & p-q & 0\\ \ (q-r)\log(d) & q-r & 0\\ \ (r-1)\log(d) + \log(f) & r & 1\end{matrix}\right|\] we would get: \[\left|\begin{matrix}\ (p-q)\log(d) & (p-q)-0 & 0\\ \ (q-r)\log(d) & (q-r) - 0 & 0\\ \ (r-1)\log(d) + \log(f) & (r)- r & 1\end{matrix}\right|\] to get: \[\left|\begin{matrix}\ (p-q)\log(d) & p-q & 0\\ \ (q-r)\log(d) & q-r & 0\\ \ (r-1)\log(d) + \log(f) & 0 & 1\end{matrix}\right|\]

OpenStudy (anonymous):

there are two ways to think about it, by expanding or by operations, whichever you prefer :)

OpenStudy (anonymous):

if we were to expand it those two elements wouldn't count because, as you said, they just get multiplied by zero

OpenStudy (anonymous):

does this help or is it still confusing? i hope i haven't made it worse lol

OpenStudy (cwrw238):

i think i'm getting it - i remember finding this stuff difficult at a-level and it never 'stuck' in my brain very well . i'll keep at it though - thanks very much - i'm glad you enjoyed doing the problem

OpenStudy (anonymous):

basically its all about using row operations. they through you a curve-ball with the logarithms and the GP

OpenStudy (anonymous):

here is a good site http://www.sosmath.com/matrix/determ1/determ1.html

OpenStudy (cwrw238):

thanks a lot

OpenStudy (anonymous):

no problem :) thanks for the juicy problem